# Persona-DB: Efficient Large Language Model Personalization for Response Prediction with Collaborative Data Refinement

### v1

@@ Abstract
The increasing demand for personalized interactions with large language models (LLMs) calls for the development of methodologies capable of accurately and efficiently identifying user opinions and preferences. Retrieval augmentation emerges as an effective strategy, as it can accommodate a vast number of users without the costs from fine-tuning. Existing research, however, has largely focused on enhancing the retrieval stage and devoted limited exploration toward optimizing the representation of the database, a crucial aspect for tasks such as personalization. In this work, we examine the problem from a novel angle, focusing on how data can be better represented for more efficient retrieval in the context of LLM customization. To tackle this challenge, we introduce Persona-DB, a simple yet effective framework consisting of a hierarchical construction process to improve generalization across task contexts and collaborative refinement to effectively bridge knowledge gaps among users. In the task of response forecasting, Persona-DB demonstrates superior efficiency in maintaining accuracy with a significantly reduced retrieval size, a critical advantage in scenarios with extensive histories or limited context windows. Our experiments also indicate a marked improvement of over 15% under cold-start scenarios, when users have extremely sparse data. Furthermore, our analysis reveals the increasing importance of collaborative knowledge as the retrieval capacity expands.
@@
Chenkai Sun, Ke Yang, Revanth Gangi Reddy, Yi R. Fung Hou Pong Chan, ChengXiang Zhai, Heng Ji University of Illinois Urbana-Champaign {chenkai5, czhai, hengji}@illinois.edu
@@ 1 Introduction
While resource-efficient, the efficacy of the method hinges on the quality of the underlying database. Current solutions for retrieval augmentation, however, have limited exploration in the facet and often depend largely on expansive yet shallow logs of user interactions and summaries, which can prevent the model from achieving optimal performance. Firstly, such an approach would make it difficult to capture the depth of user information, a vital component in the model performance in answering queries. Secondly, the approach discourages efficiency, since it needs to gather scattered information pieces to provide sufficient pertinent data for retrieval augmentation. For instance, a higher-level opinion, such as an attitude toward a political party, can often be a more broadly applicable and retrieval-efficient insight compared to an opinion about a specific entity within the party, where the former requires analyzing historical behaviors. Moreover, existing methods have yet to explore the potential synergies that could arise from intelligently leveraging the interconnectedness of different users’ data to fill knowledge gaps.
@@
Our paper delves into the critical research question: How do we address data efficiency in retrieval-augmented personalization? In other words, considering the cost of model inference, how do we maintain the accuracy of the generated response by retrieving less? To address previous gaps, we propose a solution based on the premise that the construction of more structured, self-improved, and interrelated databases can enhance the LLM’s ability to retrieve relevant information and to generalize. This leads us to the introduction of our framework Persona-DB, designed to encapsulate extensive user contexts and histories, fostering more accurate and individualized interactions.
@@
The framework is underpinned by two integral components designed to enhance user databases for more generalizable inference. The first component involves a simple yet effective stage that involves distillation and induction on the user database to encapsulate and extend user personas such as opinions and preferences. The hierarchical construction process enables the application of learned insights across diverse task contexts. The second component involves interconnecting relevant databases among users, drawing inspiration from the principles of collaborative filtering in recommendation systems. Recognizing that an individual’s data may sometimes be insufficient to address the query (e.g., due to sparse data or domain-specific gaps), we facilitate the amalgamation of other users’ data that may contain relevant interactions, based on the previous studies that users with analogous mindsets tend to exhibit similar behaviors and preferences Reimer et al. (2022); Torelli and Kaikati (2009). For instance, a user who participates in outdoor activities and values environmental sustainability, yet has only a few interactions related to renewable energy projects, can be matched with users of a similar environmental ethos who have engaged in discussions about renewable energy. This assists us in inferring the original user’s viewpoints on solar energy initiatives, leveraging shared values to supplement the absence of interactions in that domain. In our work, we specifically achieve this by matching users based on shared persona keys identified in the first component and composing information pieces from the individual’s database and the collaborative databases. Our approach aims to make databases both more precise and resource-efficient for downstream LLM usage, reducing the need to retrieve extensive data.
@@
We evaluate the method on the task of personalized response forecasting, where the downstream LLM is asked to predict users’ responses to news media. Our experimental evaluation reveals that Persona-DB consistently outperforms the baselines across various retrieval sizes, notably achieving better results when the maximal retrieval size is 10 times smaller than that of the baseline. The results indicate that our framework can maintain accuracy while requiring substantially less input data, which is particularly advantageous when dealing with extensive user histories or limited context windows in LLMs. In scenarios where the user has starkly scarce data, our framework outperforms the baseline by a substantial margin (e.g., 15% in the correlation metric), demonstrating the effectiveness of our approach to generalize under the case of new or infrequent users. In our analysis, we further explore the impact of knowledge contributed by collaborator databases, finding that its importance to overall performance escalates as the retrieval capacity grows.
@@ 2 Persona-DB
The primary challenge in enhancing retrieval augmentation for personalized LLM interactions is the effective and accurate representation of user context (i.e., the interaction data that inform the LLM’s personalization process) within the constraints of LLM context windows and amidst the vast diversity of user behaviors. To effectively represent personas, it is crucial to balance comprehensiveness and efficiency. We aim to encapsulate detailed user information, such as experiences, and opinions, while ensuring that the data is efficiently retrievable. In our context, we define personalization efficiency in terms of information density and predictive power. For instance, a personalization-efficient database will be able to maintain a downstream model with fewer retrieval items. Despite advancements in retrievers and encoders, existing methods, which predominantly rely on user logs and summaries, do not fully leverage the predictive capabilities of user personas or facilitate the exploration of relations between users. Recognizing the limitations, we introduce Persona-DB, a novel framework designed to address key challenges in efficient user context representation. Our approach is guided by two key intuitions. Firstly, high-level (or abstract) user personas, which have been demonstrated by previous social psychology works Reimer et al. (2022); Torelli and Kaikati (2009) to effectively predict human actions across different social contexts, can provide a more generalizable basis for persona representation. Secondly, inspired by collaborative filtering prevalent in recommendation systems, we explore the potential of applying similar principles in the context of LLM personalization. Specifically, our framework consists of two components, (1) the creation of a database by distilling and extending user data into more generalizable persona constructs and (2) a JOIN operation to bridge knowledge gaps between users by using keys identified in the previous stage. The workflow is depicted in Figure 1.
@@ 2.1 Hierarchical Refinement
User histories, which capture a user’s past behaviors, are often noisy, with key information dispersed and sometimes only inferable through sequential analysis. Given the scenario where an LLM (or Agent) can only use a limited amount of information due to context limitations, it becomes crucial to appropriately transform the database for more efficient downstream usage. One natural method for such transformation is creating summaries of events, yet it discourages generalization at some level. For instance, a person’s nature and beliefs (e.g., being supportive of an initiative) can typically be more generalizable across contexts than a reaction under a specific scenario (e.g., positive attitude toward a particular entity within the initiative). From previous works in social psychology, it was suggested that abstract personas such as values are generalizable to predicting future user actions under different social contexts Reimer et al. (2022). This indicates that extracting higher-level variables from user data may introduce additional benefits to personalization accuracy while requiring lighter retrieval. To achieve this step, we choose to use an instruction-tuned LLM Wei et al. (2021); Chung et al. (2022) to structure and infer user facts and opinions to automatically enhance the database. In our experiment, we use the same LLM for both hierarchical refinement and downstream prediction.
Our design schema is a hierarchy consisting of History, Distilled Persona, Induced Persona, and Cache, where an example snapshot is shown in Figure 1. The History layer contains the original user records. On the other hand, the Distilled Persona (abbreviated DP) is an induced dictionary of persona-related facts from History, such as superficial opinions and pattern analysis. The Induced Persona (abbreviated IP) layer comprises inferred higher-level information from the previous two components, offering a more abstract view of the user. For instance, DP can include specific observations, such as the user’s criticism of a governmental decision to reduce public spending on community services, while IP can include conjectures such as that the user has a broad concern regarding social justice, derived from observations in DP. Lastly, Cache is a subset of IP that contains consistent human-defined high-level persona categories to assist relevancy matching in the collaboration stage of the framework. We utilize an LLM as an analyzer to infer and enrich the database with both user facts and opinions. In particular, each layer is populated by feeding the information from the lower layers to the LLM for processing. The structure aims to create more packed user-level features to assist retrieval efficiency. By being able to utilize high-level, generalizable personas, the downstream model can apply these insights across a broad range of contexts, enhancing the model’s ability to personalize interactions.
@@ 2.2 Collaborative Refinement
While using a user’s data to predict intention is straightforward, it can introduce significant challenges when the user’s interactions are irrelevant to a domain presented in the query or when the user simply does not have many interactions. In fact, it was discovered previously that only 25% of highly active users generate 97% of the content on Twitter McClain et al. (2021), indicating that such a case is dominant in the social media context. To address the issues, we introduce a stage to interconnect knowledge between users, under the assumption that users with similar mindsets would make similar decisions Reimer et al. (2022). By enabling a user to retrieve and integrate information from relevant users, we establish a collaborative database. This enriches the user context by the inclusion of potentially more domain-matching content from collaborators, thereby enhancing the generalizability and relevance of the user context representation. As demonstrated in our experiment results, such an approach indeed brings enormous benefits in cold-start situations when the user has scarce data on the platform. Our research is also the first to probe the application of this collaborative filtering-inspired method in the relatively new domain of LLM personalization, especially in scenarios where a social network or product graph is not readily available.
@@
We then use the embedded as the key to compute similarities between the current user’s database and others, followed by retrieval and joining of the most relevant databases. Formally,
@@
Table 1: The table displays response prediction results, highlighting the best overall performance in bold. Our framework (with top-40 retrieval) consistently surpasses baseline methods, notably achieving higher accuracy than the method using full user history databases. Additionally, the results include an ablation study demonstrating the standalone effectiveness of Persona-DB without the JOIN mechanism.
@@ 3.1 Dataset Overview
To achieve personalization, a model needs to be capable of predicting users’ opinions and preferences regarding a wide range of potential responses or products. In this work, we utilize the dataset from Sun et al. (2023) (denoted as RFPN) to evaluate the model’s performance in personalized response forecasting. The dataset comprises 13.3k responses from 8.4k users to 3.8k news headlines sourced from Twitter. We utilize the dataset’s test set, in which the history length can exceed 300. We evaluate the model performance on predicting the sentiment polarity $\phi_{p}$ (categorized as Positive, Negative, or Neutral) and intensity $\phi_{int}$ (on a scale from 0 to 3), based on a given persona and news media message. Here, a persona is constructed from user attributes, including profiles and historical posts. Detailed statistics of the dataset are provided in the Appendix. The prediction task allows for reliable evaluation and analysis of how messages are perceived by different personas or social segments, simplifying the challenge of assessing text generation in personalized conversation with lengthy user contexts, which can be challenging for both LLM and Human Evaluation.
@@ 3.2 Experimental Setup
The evaluation uses Spearman and Pearson correlation coefficients, denoted as $r_{s}$ and $r$, respectively, to assess sentiment intensity predictions. The sentiment polarity classification is evaluated using Micro-F1 and Macro-F1 scores (denoted MiF1 and MaF1), which measure accuracy in multi-class scenarios. To test the efficacy of the Persona-DB components, we compare our framework against different variants that use the original user historical data as the database, including H-Retrieval, where the most relevant entries are used, and H-Recency, which uses the most recent entries. Additionally, we evaluate the performance of Persona-DB without collaborative refinement in a variant named Persona-DB w/o JOIN (abbreviated P-DB w/o J). For the experiments in 3.3 and 3.4, we use 40 as the retrieval capacity. We also include naive strategies including Random and Majority, where the former makes predictions randomly, and the latter follows the majority label. A reference to using full history, History (Full), offers insight into the performance of complete history utilization. To accommodate the context window limitation caused by an excessively large number of interactions, we limit the number of items to within the range that LLM can accept. The retrieval process utilizes an instruction-tuned LLM encoder, text-embedding-ada-002111https://platform.openai.com/docs/models/embeddings, to facilitate both the collaborative user retrieval and the downstream retrieval augmentation. Inference is performed using gpt-3.5-turbo-0613222https://platform.openai.com/docs/models/gpt-3-5-turbo via the official OpenAPI, with a fixed seed across runs. The same model is also utilized in the first stage of our framework to process user histories.
Our experiments were conducted using the PyTorch framework Paszke et al. (2019) and the Huggingface Transformers library Wolf et al. (2020). The sentiment intensity labels adhere to the definition in SemEval-2018 Task 1 Mohammad et al. (2018), incorporating both magnitude and direction in the evaluation of sentiment intensity. More details can be found in the Appendix.

### v3

@@ Abstract
The increasing demand for personalized interactions with large language models (LLMs) calls for methodologies capable of accurately and efficiently identifying user opinions and preferences. Retrieval augmentation emerges as an effective strategy, as it can accommodate a vast number of users without the costs from fine-tuning. Existing research, however, has largely focused on enhancing the retrieval stage and devoted limited exploration toward optimizing the representation of the database, a crucial aspect for tasks such as personalization. In this work, we examine the problem from a novel angle, focusing on how data can be better represented for more data-efficient retrieval in the context of LLM customization. To tackle this challenge, we introduce Persona-DB, a simple yet effective framework consisting of a hierarchical construction process to improve generalization across task contexts and collaborative refinement to effectively bridge knowledge gaps among users. In the evaluation of response prediction, Persona-DB demonstrates superior context efficiency in maintaining accuracy with a significantly reduced retrieval size, a critical advantage in scenarios with extensive histories or limited context windows. Our experiments also indicate a marked improvement of over 10% under cold-start scenarios, when users have extremely sparse data. Furthermore, our analysis reveals the increasing importance of collaborative knowledge as the retrieval capacity expands.The code is available at https://github.com/chenkaisun/Persona-DB
@@
Chenkai Sun1, Ke Yang1, Revanth Gangi Reddy1, Yi R. Fung1 Hou Pong Chan1, Kevin Small2, ChengXiang Zhai1, Heng Ji1 1University of Illinois Urbana-Champaign, 2Amazon {chenkai5, czhai, hengji}@illinois.edu
@@ 1 Introduction
While resource-efficient, the efficacy of the method hinges on the quality of the underlying database. Current solutions for retrieval augmentation, however, have limited exploration in the facet and often depend largely on expansive yet shallow logs of user interactions and summaries, which can prevent the model from achieving optimal performance. Firstly, such an approach would make it difficult to capture the depth of user information, a vital component in the model performance in answering queries. Secondly, the approach discourages retrieval efficiency, since it needs to gather scattered information pieces to provide sufficient pertinent data for retrieval augmentation. For instance, a higher-level opinion, such as an attitude toward a political party, can often be a more broadly applicable and retrieval-efficient insight compared to an opinion about a specific entity within the party, where the former requires analyzing historical behaviors. Moreover, existing methods have yet to explore the potential synergies that could arise from intelligently leveraging the interconnectedness of different users’ data to fill knowledge gaps.
@@
Our paper delves into the critical research question: How do we address context efficiency in retrieval-augmented personalization? In other words, considering the cost of model inference, how do we maintain the accuracy of the generated response by retrieving less? To address previous gaps, we propose a solution based on the premise that the construction of more structured, self-improved, and interrelated databases can enhance the LLM’s ability to retrieve relevant information and to generalize. This leads us to the introduction of our framework Persona-DB, designed to encapsulate extensive user contexts and histories, fostering more accurate and individualized interactions.
@@
The framework is underpinned by two integral components designed to enhance user databases for more generalizable inference. The first component involves a simple yet effective stage that involves distillation and induction on the user database to encapsulate and extend user personas such as opinions and preferences. The hierarchical construction process enables the application of learned insights across diverse task contexts. The second component involves interconnecting relevant databases among users, drawing inspiration from the principles of collaborative filtering in recommendation systems. Recognizing that an individual’s data may sometimes be insufficient to address the query (e.g., due to sparse data or domain-specific gaps), we facilitate the amalgamation of other users’ data that may contain relevant interactions, based on the previous studies that users with analogous mindsets tend to exhibit similar behaviors and preferences Reimer et al. (2022); Torelli and Kaikati (2009). For instance, a user who participates in outdoor activities and values environmental sustainability, yet has only a few interactions related to renewable energy projects, can be matched with users of a similar environmental ethos who have engaged in discussions about renewable energy. This assists us in inferring the original user’s viewpoints on solar energy initiatives, leveraging shared values to supplement the absence of interactions in that domain. In our work, we specifically achieve this by matching users based on shared persona keys identified in the first component and composing information pieces from the individual’s database and the collaborative databases. Our approach aims to make databases both more precise and context-efficient for downstream LLM usage, reducing the need to retrieve extensive data.
@@
We evaluate the method on Response Forecasting for News Media, where the downstream LLM is asked to predict users’ responses to news messages, and on OpinionQA, where the LLM predicts individual survey responses Sun et al. (2023); Santurkar et al. (2023). Our experimental evaluation reveals that Persona-DB consistently outperforms the baselines across various retrieval sizes, notably achieving better results when the maximal retrieval size is 10 times smaller than that of the baseline. The results indicate that our framework is advantageous when dealing with extensive user histories or limited context windows in LLMs. In scenarios where the user has starkly scarce data, our framework outperforms the baseline by a substantial margin (e.g., 11% in the correlation metric), demonstrating the effectiveness of our approach to generalize under the case of new or infrequent users. In our analysis, we further explore the impact of knowledge contributed by collaborator databases, finding that its importance to overall performance escalates as the retrieval capacity grows.
@@ 2 Persona-DB
The primary challenge in enhancing retrieval augmentation for personalized LLM interactions is the effective and accurate representation of user context (i.e., the interaction data that inform the LLM’s personalization process) within the constraints of LLM context windows and amidst the vast diversity of user behaviors. To effectively represent personas, it is crucial to balance comprehensiveness and data efficiency. We aim to encapsulate detailed user information, such as experiences and opinions, while ensuring that the data is efficiently retrievable. In our context, we define personalization efficiency in terms of information density and predictive power. For instance, a personalization-efficient database will be able to maintain a downstream model with fewer retrieval items. Despite advancements in retrievers and encoders, existing methods, which predominantly rely on user logs and summaries, do not fully leverage the predictive capabilities of user personas or facilitate the exploration of relations between users. Recognizing the limitations, we introduce Persona-DB, a novel framework designed to address key challenges in efficient user context representation. Our approach is guided by two key intuitions. Firstly, high-level (or abstract) user personas, which have been demonstrated by previous social psychology works Reimer et al. (2022); Torelli and Kaikati (2009) to effectively predict human actions across different social contexts, can provide a more generalizable basis for persona representation. Secondly, inspired by collaborative filtering prevalent in recommendation systems, we explore the potential of applying similar principles in the context of retrieval-augmented personalization. Specifically, our framework consists of two components, (1) the creation of a database by distilling and extending user data into more generalizable persona constructs and (2) a JOIN operation to bridge knowledge gaps between users by using keys identified in the previous stage. The workflow is depicted in Figure 1.
@@ 2.1 Hierarchical Refinement
User histories, which capture a user’s past behaviors, are often noisy, with key information dispersed and sometimes only inferable through sequential analysis. Given the scenario where an LLM (or Agent) can only use a limited amount of information due to context limitations, it becomes crucial to appropriately transform the database for more efficient downstream usage. One natural method for such transformation is creating summaries of events, yet it discourages generalization at some level. For instance, a person’s nature and beliefs (e.g., being supportive of an initiative) can typically be more generalizable across contexts than a reaction under a specific scenario (e.g., positive attitude toward a particular entity within the initiative). From previous works in social psychology, it was suggested that abstract personas such as values are generalizable to predicting future user actions under different social contexts Reimer et al. (2022); Ajzen (1991). This indicates that extracting higher-level variables from user data may introduce additional benefits to personalization accuracy while requiring lighter retrieval. To achieve this step, we choose to use an instruction-tuned LLM Wei et al. (2021); Chung et al. (2022) to structure and infer user facts and opinions to automatically enhance the database. In our experiment, we use the same LLM for both hierarchical refinement and downstream prediction.
Our design schema is a hierarchy consisting of History, Distilled Persona, Induced Persona, and Cache, where an example snapshot is shown in Figure 1. The History layer contains the original user records. On the other hand, the Distilled Persona (abbreviated DP) is an induced dictionary of persona-related facts from History, such as superficial opinions and pattern analysis. The Induced Persona (abbreviated IP) layer comprises inferred higher-level information from the previous two components, offering a more abstract view of the user. For instance, DP can include specific observations, such as the user’s criticism of a governmental decision to reduce public spending on community services, while IP can include conjectures such as that the user has a broad concern regarding social justice, derived from observations in DP. Lastly, Cache contains consistent human-defined high-level persona categories to assist relevancy matching in the collaboration stage of the framework. We utilize an LLM as an analyzer to infer and enrich the database with both user facts and opinions. In particular, each layer is populated by feeding the information from the lower layers to the LLM for processing. The structure aims to create more packed user-level features to assist data efficiency. By being able to utilize high-level, generalizable personas, the downstream model can apply these insights across a broad range of contexts, enhancing the model’s ability to personalize interactions.
@@ 2.2 Collaborative Refinement
While using a user’s data to predict intention is straightforward, it can introduce significant challenges when the user’s interactions are irrelevant to a domain presented in the query or when the user simply does not have many interactions. In fact, it was discovered previously that only 25% of highly active users generate 97% of the content on Twitter McClain et al. (2021), indicating that such a case is dominant in the social media context. To address the issues, we introduce a stage to interconnect knowledge between users, under the assumption that users with similar mindsets would make similar decisions Reimer et al. (2022). By enabling a user to retrieve and integrate information from relevant users, we establish a collaborative database. This enriches the user context by the inclusion of potentially more domain-matching content from collaborators, thereby enhancing the generalizability and relevance of the user context representation. As demonstrated in our experiment results, such an approach indeed brings enormous benefits in cold-start situations when the user has scarce data on the platform.
@@
We then use the embeddings as the key to compute similarities between the current user’s database and others, followed by retrieval and joining of the most relevant databases. Formally,
@@
w/o IP 43.21 43.81 60.73 47.79
w/o DP 44.42 44.51 60.44 48.72
Table 1: The table presents the results from the RFPN dataset, with the best overall performance highlighted in bold. Our framework (with top-40 retrieval) consistently surpasses baseline methods, notably achieving higher accuracy than the method using full user history databases. Additionally, the results include an ablation study demonstrating the standalone effectiveness of Persona-DB without the individual components.
@@ 3.1 Dataset Overview
To achieve personalization, a model needs to be capable of predicting users’ opinions and preferences regarding a wide range of potential responses or products. In this work, we utilize two datasets, Response Forecasting for Personas in News Media (RFPN) Sun et al. (2023) and OpinionQA Santurkar et al. (2023), to evaluate the model’s performance in personalized response prediction. Detailed statistics are provided in the Appendix.
The RFPN dataset comprises 13.3k responses from 8.4k users to 3.8k news headlines sourced from Twitter. We utilize the dataset’s test set, in which the history length can exceed 300. We evaluate the model performance on predicting the sentiment polarity $\phi_{p}$ (categorized as Positive, Negative, or Neutral) and the ordinal intensity $\phi_{int}$ (on a scale from 0 to 3), based on a given persona and news media message. Here, a persona is constructed from user attributes, including profiles and historical posts. The prediction task allows for reliable evaluation, simplifying the challenge of assessing text generation in personalized conversation with lengthy user contexts, which can be challenging for both LLM and Human. The data also naturally allows the developed model to perform personalized content moderation and recommendation. The OpinionQA dataset includes survey data from Pew Research’s American Trends Panels, and each sample consists of a single-answer multiple-choice question and the user’s selection. In OpinionQA, a persona is defined by both the respondent’s demographic information and historical responses. We specifically use Biomedical and Food Issues, Global Attitudes, and America in 2050 topics from the dataset; this allows us to assess the method’s performance across diverse scenarios.
@@ 3.2 Experimental Setup
We follow the evaluation from Sun et al. (2023) for the first task. Specifically, we assess intensity predictions using Spearman and Pearson correlation coefficients, denoted as $r_{s}$ and $r$, respectively. These metrics help quantify the association between the ordinal predicted and actual intensity scales, allowing for credits even when exact matches are not achieved. The sentiment polarity multi-class classification is evaluated using the traditional Micro-F1 and Macro-F1 scores, denoted MiF1 and MaF1, respectively. For the multiple-choice-based OpinionQA task, we measure performance using accuracy. In addition to the main results, we also perform analyses of our method on the RFPN dataset.
To test the efficacy of the Persona-DB components, we compare our framework against different variants that use the original user historical data as the database, including H-Retrieval, where the most relevant entries are used, and H-Recency, which uses the most recent entries. We also compare with  Richardson et al. (2023) (abbreviated IntSum), which uses task-aware user summaries generated by LLMs. Additionally, we evaluate the performance of Persona-DB without collaborative refinement in a variant named Persona-DB w/o JOIN (abbreviated P-DB w/o J). For the experiments in 3.3 and 3.4, we use 40 as the retrieval capacity. We also include naive strategies including Random and Majority, where the former makes predictions randomly, and the latter follows the majority label. A reference to using full history, History (Full), offers insight into the performance of complete history utilization. In the main and case studies, we set 25% as the composition ratio. To accommodate the context window limitation caused by an excessively large number of interactions, we limit the number of items to within the range that LLM can accept. The retrieval process utilizes an instruction-tuned LLM encoder, text-embedding-ada-002111https://platform.openai.com/docs/models/embeddings, to facilitate both the collaborative user retrieval and the downstream retrieval augmentation. Inference is performed using gpt-3.5-turbo-0613222https://platform.openai.com/docs/models/gpt-3-5-turbo via the official OpenAPI, with a fixed seed across runs. The same model is also utilized in the first stage of our framework to process user histories.